OpenStudy (anonymous):

If f(x)=2x, g(x)=2x-1, and h(x)=square root of x, find (f o g o h)(9) I think I have the right answer, but can someone please reassure me?

OpenStudy (imstuck):

h(9)= sqrt of 9 which is 3. Now plug 3 into g(x).

OpenStudy (imstuck):

2(3)-1= 6-1=5. Now plug 5 into f(x).

OpenStudy (imstuck):

2(5)=10. The answer is 10. Did you get that?

OpenStudy (anonymous):

@IMStuck Yes. I did. Thank you! I did it wrong the first time so I wanted to be sure.

OpenStudy (imstuck):

Good for you! God job!